A dart hits the square dartboard shown below at a random point. Find the probability that the dart lands in the shaded circular region. Each side of the dartboard is 13in , and the radius of the shaded region is 5 in.
use the value 3.14 for pi
1
Expert's answer
2020-05-21T18:05:03-0400

Probability in this happening is relation area circle with radius 5 in to square with side 13 in

p = "\pi" * R^2 / a^2 = π * 5^2 / 13^2 = π * 25/169 = 3.14 * 0.148 = 0.465

Then p = 0.465* 100 = 46.5 %